OpenStudy (anonymous):

xy=4 whats dy/dt when x=8 and dx/dt=10

OpenStudy (anonymous):

y = 4/x dy/dt = -4/x^2 times dx/dt dy/dt = -4/64 (10) dy/dt = -40/64 or -5/8

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

wait, how did u get from 4/x to -4/x^2